OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). sec2 x - 2 = tan2 x I'm not sure how to do this, could someone walk me through?

OpenStudy (loser66):

\[sec^2 x-2=tan^2 x?\]

OpenStudy (loser66):

or sec (2x) -2 = tan (2x)?

OpenStudy (anonymous):

the first one :)

OpenStudy (loser66):

no solution, friend. since 1+ tan^2 x = sec^2 x ---> 2 + tan^2 x = 1+ 1 + tan^2x = 1 + sec^2x = sec^2x leads to 0=1 ( therefore, no solution)

OpenStudy (anonymous):

oh ok i see now, thank you!